A jet heads due west at 350 km/h. It experience a 100km/h crosswind flowing in the direction N60 degreesW.Find the speed of the jet?


1
Expert's answer
2021-08-20T08:12:43-0400




"\vec v=(-350, 0), \vec u=(-100\sin(60\degree), 100\cos(60\degree))"

"\vec v+\vec u=(-350-100\sin(60\degree), 0+100\cos(60\degree))"

"speed=|\vec v+\vec u|"

"=\sqrt{(-350-100\sin(60\degree))^2+(100\cos(60\degree))^2}"

"=\sqrt{350^2+70000(\dfrac{\sqrt{3}}{2})+100^2}"

"=100\sqrt{13.25+3.5\sqrt{3}}\ (km/h)\approx439.456\ km/h"